I need help with the following: (my son asked me to help but I'm VERY rusty)
3A + B + 4C = 34
2A + 2B + C = 26
A + 2B + 3C = 23

Answer by askmemath(368)

3 equations and 3 unknowns. Take 2 equations at a time and try to eliminate 1 common variable
3A + B + 4C = 34 (1)
2A + 2B + C = 26 (2)
A + 2B + 3C = 23 (3)
Subtracting (3) fm (2)
A - 2C = 3 (4)
Multiplying (1) with 2
6A+2B+8C = 68 (5)
Subtracting (2) fm (5)
4A+7C = 42 (6)
Multiplying (4) with 4
4A-8C = 12 (7)
Subtracting (7) fm (6)
15C = 30
Dividing by 15 throughout
C = 2
Putting this back in (4)
A - 2(2) = 3
Adding 4 on both sides
A = 7
Putting this back in any of (1),(3),(2)
B = 5